Suppose you begin with 2.4 moles of sodium and 1.4 moles of fluorine. Calculate the moles of product that can be made from each reactant. 2Na+F2-----2NaF, limiting reactant and theoritical yield of NaF in grams? Please help me!!!
Answer #1

                   2Na+F2-----2NaF

                2 moles of Na react with 1 mole of F2

                2.4 moles of Na react with = 2.4/2 = 1.2 moles F2

                  Na is limiting reactant

                  F2 is excess reactant

           2 moles of Na react with F2 to form 2 moles of NaF

          2.4 moles of na react with F2 to form = 2*2.4/2 = 2.4 moles of NaF

       mass of NaF = no of moles * gram molar mass

                            = 2.4*42 = 100.8g